# Semiparametric robust estimation of population location

### v1

@@ 5.1 Simulated datasets
Let us demonstrate our practical heuristic strategy now. Consider a mixture with a dominant component consisting of $60\%$ $N(10,1)$ and a diffuse contaminating component of $40\%$ $N(20,16)$. We again generate $n=1000$ observations from this model and run Algorithm 1 on the resulting sample. The nonparametric part in the vanilla weighted EM algorithm [37] absorbed the diffuse minority and estimated the centre of the dominant population as $\hat{\theta}=19.91$, which is actually close to the centre of the contaminating component. While our practical method ran two simultaneous EM algorithms and returned $\hat{\theta}=9.98$, which is indeed very close to the true location $10.$
@@
We now consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $2$ and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$. We again generate $n=1000$ observations from this model. But now, we model the parametric part through a normal family in our working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=-0.25$, which is very close to the true centre 0, even after only approximately modeling the data through a Normal parametric family.
@@
Again, we consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $5$ (diffuse majority) and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$ (tight minority). Generate $n=1000$ observations from this model. Once again, we model the parametric part through a normal family in the working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=1.08$, which is remarkably close to the true centre 0, even after only approximately modeling the data through a Normal family.
@@ 5.2 Real datasets
This dataset records the lengths of Snapper fish, originally collected by [6]. The purpose of the data is to distinguish among the possible age groups to which an individual fish might belong—specifically, whether it comes from the current year’s spawning cohort or from that of the previous year. The empirical distribution is visibly bimodal, with modes located near 5 and 7.75 inches. The histogram shows a more pronounced peak around 5 inches, suggesting that fish from the current year’s spawning are represented in greater proportion than those from the earlier cohort.
Our algorithm estimates the average length of the dominant population to be $5.385$ inches, which is quite close to estimates calculated using more sophisticated techniques in other robust statistics literature.
@@
The second dataset is Newcomb’s classic light–speed measurements, previously analyzed by [2], [5], and others. The data records the time (in nanoseconds) required for light to traverse 7442 meters at sea level, subtracted from a baseline value of 24800 nanoseconds. The resulting histogram shows a prominent peak around 30, along with two well-known outliers at 2-2 and 44-44. We further shift the dataset by adding 44 to each observation so that all values become non-negative.
Clearly, the estimate from our algorithm is substantially close to the peak location of the larger population.
@@
This dataset contains the width–to–length ratios of $20$ beaded rectangles crafted by the Shoshoni Indians and used as decorative elements on clothing and blankets [21]. Our interest lies in estimating the central tendency of these ratios. In particular, one may ask whether the Shoshoni exhibited a preference for golden rectangles; see, for example, [22] for a discussion of golden rectangles. If such a preference existed, then the width–to–length ratios of their rectangles would be expected to cluster around the golden ratio, approximately $0.618$.
@@
The fourth dataset consists of rainfall measurements (in millimeters) recorded every fourth day during the winter months (June, July, and August) of 1981–83 in Melbourne, Australia [30]. The dataset includes a substantial outlier of approximately 300 mm. For our analysis, we use a scale-transformed version of these observations.
@@
Now let us assess the performance of our algorithm for sequential estimation of multiple significant peaks. As before, we run our algorithm on both simulated and real world datasets.
@@ 5.3 Simulated datasets
To examine a more difficult scenario, where the peaks are not very well separated and with considerable component overlap, we simulate $1000$ samples from a $3$-component gaussian mixture $0.3\>\mathcal{N}(-3,1.5)+0.4\;\mathcal{N}(0,1)+0.25\;\mathcal{N}(3,1.5).$ Our sequential extraction scheme produces 3 peaks with centres 2.866, 0.080, 2.651-2.866,\;0.080,\;2.651. Remarkably close to the true locations inspite of significant overlap between the individual components.
@@
Now let us see how our algorithm performs where there are even more components. We simulate $10000$ samples from a $5$-component gaussian mixture $0.15\;\mathcal{N}(-6,1)+0.20\;\mathcal{N}(-3,1)+0.25\;\mathcal{N}(0,1)+0.20\;\mathcal{N}(3,1)+0.20\;\mathcal{N}(6,1).$ Our algorithm produces 5 peaks with centres 6.184, 3.254, 0.013, 2.994, 6.089-6.184,\;3.254,\;0.013,\;2.994,\;6.089, which once again are sufficiently close to the true locations.
@@ 5.4 Real datasets
The classic Iris data set [12], contains 150 flowers with 50 observations from each of three species (Iris setosa, Iris versicolor, Iris virginica) and four measurements in centimetres (sepal length, sepal width, petal length, petal width); focusing on petal length, the species-wise empirical means and standard deviations are approximately $(\mu,\sigma)=(1.46,0.17)$ for setosa, $(4.26,0.47)$ for versicolor, and $(5.55,0.55)$ for virginica, so that the pooled marginal distribution of petal length is a trimodal mixture with three well-separated peaks corresponding to the three species. On applying our sequential algorithm on this dataset, we recover three distinct peaks with centers at $1.449$, $4.493$ and $5.600$, which are very good robust representatives of the empirical peak locations observed in nature, extracted from a relatively small sample.
@@
We now return to the snapper dataset. In the current-year cohort there remains a substantial fraction of fish from previous spawning seasons, which are noticeably larger in size. Although the overall population has declined due to migration and fishing pressure, a non‑negligible proportion of these older fish is still present in the sample. Applying our sequential scheme to this dataset, we identified two pronounced modes at $5.342$ and $7.499$, corresponding respectively to the newer and older spawning groups.
@@
$F^{\prime\prime}(y)=\sum_{m=0}^{M-1}c[m]\,\mathbb{K}_{h_{n}}^{\prime\prime}(y-x_{m})\quad\Rightarrow\quad\|F^{\prime\prime}\|_{\infty}\leq\sum_{m}c[m]\|\mathbb{K}_{h_{n}}^{\prime\prime}\|_{\infty}=\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{h_{n}^{3}}.$
@@
For the nonparametric part, we adopt the sieve class $\mathcal{F}_{h_{n},\Delta_{n}}=\{y\mapsto\sum_{m=0}^{M-1}c[m]K_{h_{n}}(y-x_{m}):c[m]\geq 0,\sum_{m}c[m]=1\}$ and construct the next background by smoothing the normalized nonparametric responsibilities: first map $\{\omega_{i}^{(t)}\}$ to grid coefficients $c^{(t)}[\cdot]$ via linear binning, then compute the grid convolution by zero–padded FFT, and finally use linear interpolation at $Y_{i}$ in the next E–step. This completes the EM iteration with the FFT plug–in background.
@@
Remark. Such an assumption is common in the literature, for example – [17], [14], [11], [33], [20].
@@
For each $y$, $\theta\mapsto\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}^{*}(y)\}$ is continuous on compact $\mathcal{C}$ and dominated by $H(Y)$ with $\mathbb{E}H(Y)<\infty$; see [34, Thm 19.4]. ∎
@@
Therefore, using the trivial averaging inequality $\frac{1}{n}\sum_{i=1}^{n}|a_{i}|\mathbb{I}_{\{|Y_{i}|\leq M_{n}\}}\leq\sup_{|y|\leq M_{n}}|a(y)|$ with $a(y)=\tilde{f}(y)-f_{1}^{*}(y)$, we obtain
@@
with $C_{1}$ absorbing $C_{0}$ and $|\log\pi_{0}|$. By Assumption 5, $\sup_{\theta\in\mathcal{C}}|\log f_{0}(Y;\theta)|\leq H_{0}(Y)$ and hence
@@
by dominated convergence since $H_{0},H\in L^{1}(g)$ and $M_{n}\uparrow\infty$. Also, by Assumption 4, $\Pr(|Y|>M_{n})\log(1/h_{n})\to 0.$ As $n^{-1}\sum_{i=1}^{n}\mathbb{I}_{\{|Y_{i}|>M_{n}\}}=P_{n}(|Y|>M_{n})=\Pr(|Y|>M_{n})+o_{p}(1)$, we conclude
@@
For each $y$, $\theta\mapsto\ddot{\ell}(y;\theta)$ is continuous on compact $\mathcal{C}$, and $\|\ddot{\ell}(Y;\theta)\|\leq H_{2}(Y)$ with $\mathbb{E}H_{2}(Y)<\infty$. A dominated Glivenko–Cantelli theorem for parametric classes (e.g. [34, Thm. 19.4]) yields the uniform LLN. Continuity of $\theta\mapsto\mathbb{E}[\ddot{\ell}(Y;\theta)]$ then gives the specialization at $\theta^{*}$. ∎
@@
Let $K$ be even, bounded, Lipschitz and $C^{2}$ with $\int K=1$, $\int uK(u)\,du=0$, $\int u^{2}|K(u)|\,du<\infty$ and $\|K^{\prime\prime}\|_{\infty}<\infty$. Define $K_{h}(u):=h_{n}^{-1}K(u/h_{n})$ and assume $h_{n}\downarrow 0$, $nh_{n}/\log(1/h_{n})\to\infty$. Let $\tilde{f}_{\ell,n}^{\mathrm{exact}}$ denote the exact weighted KDE with weights $\widehat{w}_{i}^{(\ell)}$:
$\tilde{f}_{\ell,n}^{\mathrm{exact}}(y):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}K_{h_{n}}(y-Y_{i}),$
and $\tilde{f}_{\ell,n}^{\mathrm{FFT}}$ its binned/FFT approximation on a grid of spacing $\Delta_{n}\downarrow 0$. Suppose there exists a deterministic window $M_{n}\uparrow\infty$ such that
@@
Fix $\eta=(\theta,\alpha)$ and abbreviate $\ell_{\eta}^{u}(y):=\ell^{(\ell)}(y;\eta,u).$ Write
@@
and that $\sup_{\eta,y}\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(y)\leq C_{0}+\log(1/h_{n})$ for all large $n$ by Assumptions 10(iv) and 11, while $\sup_{\eta}|\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)|\leq H^{(\ell)}(Y)$ by Assumption 10(vi). Using $\sum_{i}\widehat{w}_{i}^{(\ell)}\leq C_{\ell}$,
@@
Proof.
By Proposition 3, $\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}|\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta)-\mathcal{L}^{(\ell)}(\eta)|\to_{p}0.$ By Assumption 10(v), $L^{(\ell)}$ is continuous on the compact set $\mathcal{C}\times\Pi_{\min}$ with a unique maximiser $\eta_{\ell}^{*}$. Wald’s argmax theorem yields $\hat{\eta}_{\ell,n}\to_{p}\eta_{\ell}^{*}$. ∎
@@
Assumption 15 (Oracle differentiability and information at stage $\ell$).
@@
$J^{(\ell)}(\eta_{\ell}^{*}):=\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})^{\top}\big]=-\,\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}^{2}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\big]$
@@
$\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})\big).$
Assumption 17 (Oracle–equivalence of plug–in score/Hessian).
@@
By Assumption 16, $\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\rightsquigarrow\mathcal{N}(0,J^{(\ell)}(\eta_{\ell}^{*}))$. By Assumption 17, $\sqrt{n}(\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})-S_{n}^{(\ell)}(\eta_{\ell}^{*}))\to_{p}0$ and $\sup_{\eta}\|\tilde{H}_{n}^{(\ell)}(\eta)-H_{n}^{(\ell)}(\eta)\|\to_{p}0.$ Assumption 15 implies $H_{n}^{(\ell)}(\eta)\to_{p}-J^{(\ell)}(\eta)$ uniformly on $\mathcal{C}\times\Pi_{\min}$, so $\tilde{H}_{n}^{(\ell)}(\tilde{\eta}_{\ell,n})\to_{p}-J^{(\ell)}(\eta_{\ell}^{*})$. By Slutsky’s theorem,
$\sqrt{n}(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$

### v2

@@ 5.1 Simulated datasets
Let us demonstrate our practical heuristic strategy now. Consider a mixture with a dominant component consisting of $60\%$ $N(10,1)$ and a diffuse contaminating component of $40\%$ $N(20,16)$. We again generate $n=1000$ observations from this model and run Algorithm 1 on the resulting sample. The nonparametric part in the vanilla weighted EM algorithm [You2024sequential] absorbed the diffuse minority and estimated the centre of the dominant population as $\hat{\theta}=19.91$, which is actually close to the centre of the contaminating component. While our practical method ran two simultaneous EM algorithms and returned $\hat{\theta}=9.98$, which is indeed very close to the true location $10.$
@@
Consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $2$ and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$. We generate $n=500$ samples from this model. But now, we model the parametric part through a normal family in our working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=-0.25$, which is very close to the true centre 0, even after only approximately modeling the dominant component through a Normal parametric family.
@@
Again, we consider a mixture with a dominant component consisting of $60\%$ Cauchy with centre 0, scale $5$ (diffuse majority) and a contaminating component of $40\%$ Cauchy with centre $10$, scale $1$ (tight minority). Generate $n=1000$ observations from this model. Once again, we model the dominant component through a normal family in the working log-likelihood and run Algorithm 1 on the sample. The estimated centre of the dominant population is $\hat{\mu}=1.08$, which is remarkably close to the true centre 0, even when we only approximately model the larger component through a Normal family.
@@ 5.2 Real datasets
This dataset records the lengths of Snapper fish, originally collected by [cassie1954some]. The purpose of the data is to distinguish among the possible age groups to which an individual fish might belong—specifically, whether it comes from the current year’s spawning cohort or from that of the previous year. The empirical distribution is visibly bimodal, with modes located near 5 and 7.75 inches. The histogram shows a more pronounced peak around 5 inches, suggesting that fish from the current year’s spawning are represented in greater proportion than those from the earlier cohort.
Our algorithm estimates the average length of the dominant population to be $5.385$ inches, which is very close to estimates calculated using more sophisticated techniques in other robust statistics literature.
@@
The second dataset is Newcomb’s classic light–speed measurements, previously analyzed by [Basu1998], [brown1993approximate], and others. The data records the time (in nanoseconds) required for light to traverse 7442 meters at sea level, subtracted from a baseline value of 24800 nanoseconds. The resulting histogram shows a prominent peak around 30, along with two well-known outliers at 2-2 and 44-44. We further shift the dataset by adding 44 to each observation so that all values become non-negative. Clearly, the estimate from our algorithm is substantially close to the peak location of the larger population and coincides with other estimates from more sophisticated robust techniques.
@@
The next dataset contains the width–to–length ratios of $20$ beaded rectangles crafted by the Shoshoni Indians and used as decorative elements on clothing and blankets [larsen2005introduction]. Our interest lies in estimating the central tendency of these ratios. In particular, one may ask whether the Shoshoni exhibited a preference for golden rectangles; see, for example, [Livio2002GoldenRatio] for a discussion of golden rectangles. If such a preference existed, then the width–to–length ratios of their rectangles would be expected to cluster around the golden ratio, approximately $0.618$.
@@
The fourth dataset consists of rainfall measurements (in millimeters) recorded every fourth day during the winter months (June, July, and August) of 1981–83 in Melbourne, Australia [staudte2011robust]. The dataset includes a substantial outlier of approximately 300 mm. For our analysis, we use a scale-transformed version of these observations.
@@
Now let us assess the performance of our algorithm for sequential extraction of multiple significant peaks. As before, we run our algorithm on both simulated and real world datasets.
@@ 5.3 Simulated datasets
To examine a more complicated scenario, where the peaks are not very well separated and with considerable overlap between components, we simulate $1000$ samples from a $3$-component gaussian mixture $0.3\>\mathcal{N}(-3,1.5)+0.4\;\mathcal{N}(0,1)+0.25\;\mathcal{N}(3,1.5).$ Our sequential extraction scheme produces 3 peaks with centres 2.866, 0.080, 2.651-2.866,\;0.080,\;2.651. Remarkably close to the true locations inspite of significant overlap between the individual components.
@@
Now let us see how our algorithm performs where there are even more components. We simulate $10000$ samples from a $5$-component gaussian mixture $0.15\;\mathcal{N}(-6,1)+0.20\;\mathcal{N}(-3,1)+0.25\;\mathcal{N}(0,1)+0.20\;\mathcal{N}(3,1)+0.20\;\mathcal{N}(6,1).$ Our algorithm produces 5 peaks with centres 6.184, 3.254, 0.013, 2.994, 6.089-6.184,\;3.254,\;0.013,\;2.994,\;6.089, which once again are reasonably close to the true locations.
@@ 5.4 Real datasets
The classic Iris data set [fisher1936iris], contains 150 flowers with 50 observations from each of three species (Iris setosa, Iris versicolor, Iris virginica) and four measurements in centimetres (sepal length, sepal width, petal length, petal width); focusing on petal length, the species-wise empirical means and standard deviations are approximately $(\mu,\sigma)=(1.46,0.17)$ for setosa, $(4.26,0.47)$ for versicolor, and $(5.55,0.55)$ for virginica, so that the pooled marginal distribution of petal length is a trimodal mixture with three well-separated peaks corresponding to the three species. On applying our sequential algorithm on this dataset, we recover three distinct peaks with centers at $1.449$, $4.493$ and $5.600$, which suggest the existence of three distinct species in the dataset. These estimates are also quite accurate and robust representatives of the empirical means observed in nature even though we have estimated these from a relatively small sample.
@@
We now return to the snapper dataset. In the current-year cohort there remains a significant fraction of fish from previous spawning seasons, which are noticeably larger in size. Although the overall population has declined due to migration and fishing practices, a non‑negligible proportion of these older fish is still present in the sample. Applying our sequential scheme to this dataset, we identified two pronounced modes at $5.342$ and $7.499$, corresponding respectively to the newer and older spawning groups.
@@
$F^{\prime\prime}(y)=\sum_{m=0}^{M-1}c[m]\,\mathbb{K}_{h_{n}}^{\prime\prime}(y-x_{m})\quad\Rightarrow\quad\|F^{\prime\prime}\|_{\infty}\leq\sum_{m}c[m]\|\mathbb{\mathbb{K}}_{h_{n}}^{\prime\prime}\|_{\infty}=\frac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{h_{n}^{3}}.$
@@
For the nonparametric part, we adopt the sieve class $\mathcal{F}_{h_{n},\Delta_{n}}=\{y\mapsto\sum_{m=0}^{M-1}c[m]\mathbb{K}_{h_{n}}(y-x_{m}):c[m]\geq 0,\sum_{m}c[m]=1\}$ and construct the next background by smoothing the normalized nonparametric responsibilities: first map $\{\omega_{i}^{(t)}\}$ to grid coefficients $c^{(t)}[\cdot]$ via linear binning, then compute the grid convolution by zero–padded FFT, and finally use linear interpolation at $Y_{i}$ in the next E–step. This completes the EM iteration with the FFT plug–in background.
@@
Remark. Such an assumption is common in the literature, for example – [Hansen2008], [GineGuillou2002], [EinmahlMason2005], [Tsybakov2009], [Jiang2017].
@@
For each $y$, $\theta\mapsto\log\{\pi_{0}f_{0}(y;\theta)+\pi_{1}f_{1}^{*}(y)\}$ is continuous on compact $\mathcal{C}$ and dominated by $H(Y)$ with $\mathbb{E}H(Y)<\infty$; see [vaart1998, Thm 19.4]. ∎
@@
Therefore, using the trivial averaging inequality $\frac{1}{n}\sum_{i=1}^{n}|a_{i}|\mathbb{I}_{\{|Y_{i}|\leq M_{n}\}}\leq\sup_{|y|\leq M_{n}}|a(y)|$
with $a(y)=\tilde{f}(y)-f_{1}^{*}(y)$, we obtain
@@
with $C_{1}$ absorbing $C_{0}$ and $|\log\pi_{0}|$.
By Assumption 5, $\sup_{\theta\in\mathcal{C}}|\log f_{0}(Y;\theta)|\leq H_{0}(Y)$
and hence
@@
by dominated convergence since $H_{0},H\in L^{1}(g)$ and $M_{n}\uparrow\infty$.
Also, by Assumption 4, $\Pr(|Y|>M_{n})\log(1/h_{n})\to 0.$ As $n^{-1}\sum_{i=1}^{n}\mathbb{I}_{\{|Y_{i}|>M_{n}\}}=P_{n}(|Y|>M_{n})=\Pr(|Y|>M_{n})+o_{p}(1)$, we conclude
@@
For each $y$, $\theta\mapsto\ddot{\ell}(y;\theta)$ is continuous on compact $\mathcal{C}$, and $\|\ddot{\ell}(Y;\theta)\|\leq H_{2}(Y)$ with $\mathbb{E}H_{2}(Y)<\infty$. A dominated Glivenko–Cantelli theorem for parametric classes (e.g. [vaart1998, Thm. 19.4]) yields the uniform LLN. Continuity of $\theta\mapsto\mathbb{E}[\ddot{\ell}(Y;\theta)]$ then gives the specialization at $\theta^{*}$. ∎
@@
Let $\mathbb{K}$ be even, bounded, Lipschitz and $C^{2}$ with $\int\mathbb{K}=1$, $\int u\mathbb{K}(u)\,du=0$, $\int u^{2}|\mathbb{K}(u)|\,du<\infty$ and $\|\mathbb{K}^{\prime\prime}\|_{\infty}<\infty$. Define $\mathbb{K}_{h_{n}}(u):=h_{n}^{-1}\mathbb{K}(u/h_{n})$ and assume $h_{n}\downarrow 0$, $nh_{n}/\log(1/h_{n})\to\infty$. Let $\tilde{f}_{\ell,n}^{\mathrm{exact}}$ denote the exact weighted KDE with weights $\widehat{w}_{i}^{(\ell)}$:
$\tilde{f}_{\ell,n}^{\mathrm{exact}}(y):=\sum_{i=1}^{n}\widehat{w}_{i}^{(\ell)}\mathbb{K}_{h_{n}}(y-Y_{i}),$
and $\tilde{f}_{\ell,n}^{\mathrm{FFT}}$ its FFT approximation on a grid of spacing $\Delta_{n}\downarrow 0$. Suppose there exists a deterministic window $M_{n}\uparrow\infty$ such that
@@
Fix $\eta=(\theta,\alpha)$ and abbreviate $\ell_{\eta}^{u}(y):=\ell^{(\ell)}(y;\eta,u).$
Write
@@
and that $\sup_{\eta,y}\ell_{\eta}^{\tilde{f}_{\ell,n}^{\mathrm{FFT}}}(y)\leq C_{0}+\log(1/h_{n})$ for all large $n$ by Assumptions 10(iv) and $1$, while $\sup_{\eta}|\ell_{\eta}^{f_{\mathrm{bg}}^{(\ell),*}}(Y)|\leq H^{(\ell)}(Y)$ by Assumption 10(vi). Using $\sum_{i}\widehat{w}_{i}^{(\ell)}\leq C_{\ell}$,
@@
By Proposition 3, $\sup_{\eta\in\mathcal{C}\times\Pi_{\min}}|\tilde{\mathcal{L}}_{n}^{(\ell)}(\eta)-\mathcal{L}^{(\ell)}(\eta)|\to_{p}0.$ By Assumption 10(v), $L^{(\ell)}$ is continuous on the compact set $\mathcal{C}\times\Pi_{\min}$ with a unique maximiser $\eta_{\ell}^{*}$. Wald’s argmax theorem yields $\hat{\eta}_{\ell,n}\to_{p}\eta_{\ell}^{*}$.
@@
Assumption 15.
@@
$I^{(\ell)}(\eta_{\ell}^{*}):=\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\nabla_{\eta}\ell^{(\ell)}(Y;\eta_{\ell}^{*})^{\top}\big]=-\,\mathbb{E}_{f^{(\ell)}}\big[\nabla_{\eta}^{2}\ell^{(\ell)}(Y;\eta_{\ell}^{*})\big]$
@@
$\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,I^{(\ell)}(\eta_{\ell}^{*})\big).$
Assumption 17.
@@
By Assumption 16, $\sqrt{n}\,S_{n}^{(\ell)}(\eta_{\ell}^{*})\rightsquigarrow\mathcal{N}(0,I^{(\ell)}(\eta_{\ell}^{*}))$.
By Assumption 17, $\sqrt{n}(\tilde{S}_{n}^{(\ell)}(\eta_{\ell}^{*})-S_{n}^{(\ell)}(\eta_{\ell}^{*}))\to_{p}0$ and $\sup_{\eta}\|\tilde{H}_{n}^{(\ell)}(\eta)-H_{n}^{(\ell)}(\eta)\|\to_{p}0.$ Assumption 15 implies $H_{n}^{(\ell)}(\eta)\to_{p}-I^{(\ell)}(\eta)$ uniformly on $\mathcal{C}\times\Pi_{\min}$, so $\tilde{H}_{n}^{(\ell)}(\tilde{\eta}_{\ell,n})\to_{p}-I^{(\ell)}(\eta_{\ell}^{*})$. By Slutsky’s theorem,
$\sqrt{n}(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*})\ \rightsquigarrow\ \mathcal{N}\big(0,I^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$